OpenStudy (anonymous):

If sin θ = 1/3 and tan θ < 0, what is the value of cos θ?

OpenStudy (anonymous):

draw a triangle

OpenStudy (anonymous):



OpenStudy (anonymous):

there is a picture of an angle whose sine is \(\frac{1}{3}\) find the missing side via pythagoras

OpenStudy (anonymous):

you can pretty much do it in your head what is \(3^2\)?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

and \(1^2\)?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

and \(9-1\)?

OpenStudy (anonymous):

8

OpenStudy (anonymous):

so answer for the missing side is \(\sqrt{8}\) now two more steps

OpenStudy (anonymous):

2√2

OpenStudy (anonymous):



OpenStudy (anonymous):

yeah that too now the cosine is adjacent over hypotenuse, but don't forget that \(\tan(\theta)<0\) which means the cosine must be negative

OpenStudy (anonymous):

-√8/3

OpenStudy (anonymous):

or -(2√2)/3

OpenStudy (anonymous):

yes probably your teacher will like the second one more

OpenStudy (anonymous):

YES! thank you! you're gr8 :)

OpenStudy (anonymous):

lol yw